OpenStudy (ammarah):

What term is needed to add to each side to complete the square... 4x^ +12x =16

OpenStudy (ammarah):

do i divide by two on the left?

OpenStudy (ammarah):

i got 9 as an anwer is that correct??

OpenStudy (ammarah):

@geerky42

geerky42 (geerky42):

Divide both sides by 4

OpenStudy (ammarah):

wait is my answer correct?

OpenStudy (ammarah):

@geerky42

OpenStudy (anonymous):

First, divide both sides by 4, then take 1/2 of the coefficient of the x term, and square it. What do you get?

OpenStudy (ammarah):

9...

OpenStudy (anonymous):

Not 9.

OpenStudy (jdoe0001):

\(\bf 4x^2 +12x =16 \quad ?\)

OpenStudy (ammarah):

wait what do u mean by divide bothe sides? can u please show me step by step...

OpenStudy (ammarah):

yes @jdoe0001

OpenStudy (anonymous):

Divide both sides by 4, you get an x^2 + 3x on the left side. Take 1/2 of 3 which is 3/2. Now square 3/2, you get 9/4. So you add 9/4 to both sides!

OpenStudy (jdoe0001):

\(\bf 4x^2 +12x =16\qquad \textit{let's group and common factor}\\ 4(x^2 +3x) =16\qquad thus\\ 4(x^2 +3x+\square^2) =16\)

OpenStudy (ammarah):

ohhk how about this one 4x@ -20x =25 i got 25/4 is that correct>

OpenStudy (ammarah):

im sorry the at sign is supposed to be squared

OpenStudy (anonymous):

You are correct, add 25/4.

OpenStudy (ammarah):

okay i have another wuestion.... it says to solve each eqation by completing the square. x^ +6x+3=0

OpenStudy (anonymous):

So first subtract 3 from both sides, so you get x^2 + 6x = -3 Now add 9 to both sides, so you get x^2 + 6x + 9 = - 3 + 9 (x+3)^2 = 6 (x + 3) = +/- sqrt 6 Hence, x = -3 +/- sqrt (6)

OpenStudy (ammarah):

i dont get this question: its the same directions as the previous one... x^ -3x-7=0

OpenStudy (anonymous):

Add 7 to both sides. You now get x^2 - 3x = 7 Add 9/4 to both sides, you get x^2 -3x + 9/4 = 7 + 9/4 (x - 3/2)^2 = 37/4 So, (x - 3/2) = +/- sqrt (37/4) Hence, x = 3/2 +/- sqrt (37/4)

OpenStudy (ammarah):

thaniks so much i have a couple of more questions.............im gonna post it on another one..